How many moles of allicin are present in 6.00 mg of this substance?
1
Expert's answer
2019-09-26T05:53:19-0400


Molecular formula is C6H10S2O (molar mass 162.27 g/mol)


"n=\frac{m}{M} = \frac{6.00\times 10^{-3}g}{162.27\frac{g}{mol}}=3.70\times10^{-5}mol"